A dealer makes up a 60-lb mixture of almonds and cashews. if the almonds cost $4.70 per lb and the cashews cost $5.50 per lb, how many lbs of each must be used in order for the mixture to cost $4.90 per lb?
I tried this
value + value = 4.90
4.70x + 5.50(60-x)=4.90
4.70x + 330 - 5.50x= 4.90
-.8x+330-330=4.90-330
-.8x=-325.1
x=406.375
I am completely confused . Am i even on the right track?
HELP PLEASE

Answer by ptaylor(1894)

YOU ARE TOTALLY ON THE RIGHT TRACK!!!! ---JUST A SLIGHT OVERSIGHT BETWEEN YOUR WORK AND THE CORRECT ANSWER. GOOD WORK!!!! I THINK THAT YOU CAN NOW FINISH THE WORK THAT I STARTED (BELOW)
Value + Value = Value (of the final mixture)
Let x= lbs of Almonds
Then 60-x=lbs of Cashews
Now we know that the value of the Almonds(4.70x) plus the value of the Cashews (5.50(60-x) equals the value of the final mixture 4.90(60).
So our equation to solve is:
4.70x+5.50(60-x)=4.90(60)
I think that you are totally capable of finishing this.

Hope this helps------ptaylor